Really short question! Please help me to solve, thank you!
(30%)Q2 (Poisson regression): We collected n 15 independent count observations {Vi : 1,..., 15 and their corresponding covariates (i 1,..., 15). Assume the relationship between Vi and xỉ (for i-: 1, , 15) is yi ~ Poisson(A) and log(A) α+ßxi. Please 1) write down the likelihood function L(a, B|x, y) of the Poisson regression model; 2) derive the Newton method for maxmizing L(a, BIx, y); 3) implement the Newton method using R to get MLE of (α, β) i -0.30 0.32 0.41 0.62 -0.21 0.31 0.41 0.81 i 2 2 10 1 22 0 6 9 34 0.50 -0.21 -0.20 0.70 0.10 0.13 0.69 5 0 21 3 2 29
Answer #1

Here given that dependent (response) variable is Y, and we collected data n=15 with their corresponding independent variable X. Given that,

yi ~ Poisson (lambda _{i}) ...... i = 1, 2, ........... 15

And Log(lambda _{i})=alpha + eta x_{i}

1) Let Likelihood function of the Poisson Regression model is as

Lla, β.yi) yi

= prod_{i=1}^{n} rac {e^{-e^{alpha + eta x_{i}}} [e^{alpha+eta x_{i}}]^{y_{i}}}{y_{i}!}

The Log Likelihood function is as

l(alpha, eta ; y_{i})= -sum_{i=1}^{n}e^{alpha + eta x_{i}}+ sum_{i=1}^{n} y_{i}(alpha+eta x_{i}) - sum_{i=1}^{n}log(y_{i}!)

2) The first order condition for a maximum is

rac{d}{dlambda}l(lambda; x_{1},......, x_{n}) = 0

The first derivative of the log-likelihood with respect to the parameter lambda is

rac{d}{dlambda}l(lambda; x_{1},......, x_{n}) = rac{d}{dlambda} (-nlambda - sum_{j=1}^{n}ln(x_{j}!) + ln(lambda)sum_{j=1}^{n}x_{j})

=-n + rac{1}{lambda}sum_{j=1}^{n}x_{j}

Impose that the first derivative be equal to zero, and get

lambda =rac{1}{n}sum_{j=1}^{n}x_{j}

3) By using R get MLE as

R RGui (64-bit) File Edit View Misc Packages Windows Help R R Console citation ) on how to cite R or R packages in publications. Type demo () for some demos, help )for on-line help, or help.start )for an HTML browser interface to help. Type to quit R. [Previously saved workspace restored] x <-c(-0.30, 0.32, 0.41, 0.62, -0.21, 0.31, 0.41, 0.81, 0.50, -0.21, -0.20, $ > y<-c(2, 10, 11, 22, 0, 6, 9, 34, 5, 0, 1, 21, 3, 2, 29) >p_data <- data.frame (y,x) >p glm <- glm (yx, family-poisson, data-p data) >summary (p glm) Call: glm (formula = y ~ x, family = poisson, data -p data) Deviance Residuals: Min 10 Mediar 30 Max -2.31138 -0.74186 -0.01253 0.75646 1.31923 Coefficients: Intercept) -7209 3.5419 Estimate Std. Error z value Pr (>Iz|) 0.2264 3.183 0.00146 0.3590 9.866 < 2e-16 Signif. codes: 0*0.001 **O.01 0.05. 0.11 (Dispersion parameter for poisson family taken to be 1) Null deviance 163.093 on 14 degrees of freedom Residual deviance: 16.065 on 13 degrees of freedom AIC: 70.291 Number of Fisher Scoring iterations: 5

Here we get the value as

alpha = 0.7209 and eta= 3.5419

So the model is as

y = 0.7209 + 3.5419*x